OpenStudy (anonymous):

Is this equation separable? \[u_{xx}+u_{xt}+u_t=0\] I simplify down to \[X''(x)T(t)+X'(x)T'(t)+X(x)T'(t)=0\] Which does not look separable to me.

OpenStudy (anonymous):

Even if you divided by X(x)T(t) \[\frac{X''(x)}{X(x)}+\frac{X'(x)T'(t)}{X(x)T(t)}+\frac{T'(t)}{T(t)}=0\] Which is almost separable but I don't think you can do anything with the middle term.

OpenStudy (anonymous):

Turns out it is: \[X''(x)T(t)+X'(x)T'(t)+X(x)T'(t)=0\] \[X''(x)T(t)+T'(t)[X'(x)+X(x)]=0\] Divide by T(t)(X'(x)+X(x)) \[\frac{X''(x)}{X'(x)+X(x)}+\frac{T'(t)}{T(t)}=0\] \[\frac{X''(x)}{X'(x)+X(x)}=-\frac{T'(t)}{T(t)}\] \[\rightarrow \frac{X''(x)}{X'(x)+X(x)}=-\lambda , \frac{T'(t)}{T(t)}=-\lambda \]